1.The order is chloral hydrate 200 mg. The bottle reads chloral hydrate 0.1 g/capsule. How many capsules do you give?

2.The nurse is preparing to administer IV fluids to the 13 kg child who has dehydration. The daily IV fluid requirement is to administer 1000 ml + 50 ml/kg over 10 kg. How many milliliters should the nurse calculate to administer the IV therapy correctly?
Answer #1

ANSWER.

1) Calculation of number of capsules required.

            Ordered dose = 200 mg

             Available dose = 0.1g per capsule.

     Convert g to mg for calculation

              1 g = 1000 mg

      So

           Available dose = 0.1 * 1000

                                   = 100 mg per capsule.

Number of capsules required = Ordered dose/Available dose

                                                  = 200/100

                                                  = 2 capsules.

2) Calculation of ml of drug required.

          Weight of the child = 13 kg.

       Daily fluid requirement of a child over 10 kg = 1000 ml + 50 ml/kg

So

      ml of drug reuired for this child = 1000 + 50 * 13

                                                        = 1000 + 650

                                                        = 1650 ml.